Hi, I am stuck on the following question and can't seem to find a solution:
23% of a sum of money is $142.60; what is 20% of it?
Possible outcomes A) $120.08 B) $62 C) $124 D) $95.10
Thanks for your help!
Gladys

Found 2 solutions by ewatrrr, jorel1380:
Answer by ewatrrr(24785)

 
Hi,best to think of it as a 2-part question
Let x represent that sum of money
.23x = $142.60
x = $142.60/.23 = $620, that sum of money
AND
.20*$620 = $124

Answer by jorel1380(3719)

23/20=142.60/x
2852=23x
2852/23=x
x=$124.00..